OpenStudy (anonymous):

5. You want to divide a square piece of paper into two equivalent triangles. If the square measures 20 cm on each side, what will the third side of each triangle measure? (1 point) 8.9 cm 20 cm 28.3 cm 40 cm

OpenStudy (mathstudent55):

 In a right triangle you have a^2 + b^2 = c^2, but since a and b are sides of a square, a = b, so you have: a^2 + a^2 = c^2 2a^2 = c^2 c^2 = 2a^2 c = a*sqrt(2) In this case, a = 20, so c = a*sqrt(2) = 20*sqrt(2) = 28.28427...

OpenStudy (anonymous):

so im guessing 28.3

OpenStudy (mathstudent55):

Yes, that is 28.28427... rounded of to the nearest tenth.